# End to End Collaborative Synthetic Data Generation

### v1

@@ 4.3 Secure Evaluation of Synthetic Data
1:Secret shares of binned data $[\![D]\!]$, Secret shares of binned synthetic data $[\![\hat{D}]\!]$, number of patients $N$, Domain of gene features $\Omega_{g}={0,1,2,3}$ and classes $\Omega_{y}={0,1,2,3,4}$
2:Secret shares of workload error $[\![e]\!]$
3:Initialize $[\![e]\!]$
4:$[\![\mu_{g}]\!]$ , $[\![\mu_{y}]\!]$ , $[\![\mu_{g,y}]\!]\leftarrow$ Run Lines 1–32 from Protocol 4 on $[\![D]\!]$
5:for every value $[\![x]\!]$ in $[\![rc_{g}]\!]$, $[\![rc_{l}]\!]$, $[\![rc_{l}g]\!]$ do $\triangleright$ Adding noise to guarantee DP
6:     $[\![x]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],(1/N))$
@@
8:$[\![\hat{\mu}_{g}]\!]$ , $[\![\hat{\mu}_{y}]\!]$ , $[\![\hat{\mu}_{g,y}]\!]\leftarrow$ Run Lines 1–32 from Protocol 4 on $[\![\hat{D}]\!]$
9:for every value $[\![x]\!]$ in $[\![\hat{\mu}_{g}]\!]$, $[\![\hat{\mu}_{l}]\!]$, $[\![\hat{\mu}_{l}g]\!]$ do
10:     $[\![x]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],(1/N))$
@@
12:for (every value $[\![\hat{x}]\!]$ in $[\![\hat{\mu}_{g}]\!]$, $[\![\hat{\mu}_{l}]\!]$, $[\![\hat{\mu}_{l}g]\!]$) and ($[\![x]\!]$ in $[\![{\mu}_{g}]\!]$, $[\![{\mu}_{l}]\!]$, $[\![{\mu}_{l}g]\!]$)  do
13:     $[\![e]\!]\leftarrow[\![e]\!]+\pi_{\mathsf{ABS}}([\![x]\!]-[\![\hat{x}]\!])$
@@
We note that Line 21 in Algorithm 1 decrements $[\![c]\!]$ for a data custodian only if both the metrics meet the given thresholds. The data custodians will secret share an array of thresholds for each metric to this end.
@@ 4.4 Empirical Evaluation
Table 1: Quality evaluation of the generated synthetic genomic data. We run the framework with $|S|=3$. We set $k=5$ and tune the number of iterations of the Private-PGM model. Once the final synthetic data is published, we compute WLE between synthetic and combined real data and report difference of machine learning utility scores (accuracy and F1 scores with logistic regression $\Delta$LR and decision tree $\Delta$DT) between synthetic and combined real data (note: this evaluation is not part of the framework). We report averages taken over 3 runs.
WLE |$\Delta$LR Accuracy| |$\Delta$LR F1| |$\Delta$DT Accuracy| |$\Delta$DT F1|
@@
Performance evaluation. We run our framework for different threat models [35, 36, 37] and report the benchmarking runtimes for running the MPC protocols for leukemia data in Table 2. We run for the similar setup as mentioned above, and therefore the input data size $N=945$. For 3PC passive, the parts of the framework with MPC protocol run for a total of $\sim 2.5$ hours – one run of $\pi_{\mathsf{BIN}}$ takes 82.93 secs., $\pi_{\mathsf{INV-BIN}}$ takes 11.83 secs., $\pi_{\mathsf{NOISY-MARG}}$ takes 249.50 secs. ,$\pi_{\mathsf{LR}}$ with 150 epochs takes 15.42 secs. and $\pi_{\mathsf{WLE}}$ takes 158.66 secs. Table 2 reports performance evalaution for different threat models. The observed runtimes are in-line with the literature and demonstrate that an MPC based framework is feasible for end-to-end synthetic data generation where privacy is of utmost importance compared to the computational time required to generate valuable synthetic data. Future work will focus on optimizing these protocols and proposing new optimized MPC protocols for the proposed framework.
@@ Appendix A Secure Multi-Party Computation
Replicated sharing-based 3PC. To give the reader an understanding of how MPC protocols work, we give a brief description of a specific MPC protocol based on replicated secret sharing with three computing parties [35]. A private value $x$ in $\mathbb{Z}_{q}$ is secret shared among servers (parties) $S_{1},S_{2},$ and $S_{3}$ by selecting uniformly random shares $x_{1},x_{2},x_{3}\in\mathbb{Z}_{q}$ such that $x_{1}+x_{2}+x_{3}=x\mod{q}$, and giving $(x_{1},x_{2})$ to $S_{1}$, $(x_{2},x_{3})$ to $S_{2}$, and $(x_{3},x_{1})$ to $S_{3}$. We denote a secret sharing of $x$ by $[\![x]\!]$.
@@
To multiply secret shared values $[\![x]\!]$ and $[\![y]\!]$, we can use the fact that $x\cdot y=(x_{1}+x_{2}+x_{3})(y_{1}+y_{2}+y_{3})\mod{q}$. This means that $S_{1}$ computes $z_{1}=x_{1}\cdot y_{1}+x_{1}\cdot y_{2}+x_{2}\cdot y_{1}\mod{q}$, $S_{2}$ computes $z_{2}=x_{2}\cdot y_{2}+x_{2}\cdot y_{3}+x_{3}\cdot y_{2}\mod{q}$, and $S_{3}$ computes $z_{3}=x_{3}\cdot y_{3}+x_{3}\cdot y_{1}+x_{1}\cdot y_{3}\mod{q}$. The next step is for the servers to obtain an additive secret sharing of 00 by choosing random values $u_{1},u_{2},u_{3}$ such that $u_{1}+u_{2}+u_{3}=0\mod{q}$. This can be done using pseudorandom functions. Each server $S_{i}$ then computes $v_{i}=z_{i}+u_{i}\mod{q}$. Finally, $S_{1}$ sends $v_{1}$ to $S_{3}$, $S_{2}$ sends $v_{2}$ to $S_{1}$, and $S_{3}$ sends $v_{3}$ to $S_{2}$. This allows the servers $S_{1},S_{2}$, and $S_{3}$ to obtain the replicated secret shares $(v_{1},v_{2})$, $(v_{2},v_{3})$, and $(v_{3},v_{1})$, respectively, of the value $v=x\cdot y$.
@@ A.2 MPC Primitives
  • Secure multiplication $\pi_{\mathsf{MUL}}$ : At the start of this protocol, the parties have secret sharings $[\![a]\!]$ and $[\![b]\!]$; at the end, then they have a secret share of $c=a.b$.
  • Secure equality test $\pi_{\mathsf{EQ}}$ : At the start of this protocol, the parties have secret sharings $[\![a]\!]$; at the end if $a=0$, then they have a secret share of $1$, else a secret sharing of 00.
  • Secure less than test $\pi_{\mathsf{LT}}$ : At the start of this protocol, the parties have secret sharings $[\![a]\!]$ and $[\![b]\!]$ of integers $a$ and $a$; at the end of the protocol they have a secret sharing of $1$ if $a<b$, and a secret sharing of 00 otherwise.
  • Secure less than test $\pi_{\mathsf{ABS}}$ : At the start of this protocol, the parties have secret sharings $[\![a]\!]$; at the end of the protocol they have a secret sharing of of its absolute value.
@@ A.3 Protocol for Gaussian Noise
1:Secret shares of drawn sample $[\![\gamma]\!]$
2:Initialize vector $[\![\gamma]\!]$ of length $\max(\omega_{r})$ with 00s
@@
6:         $[\![sum]\!]\leftarrow[\![sum]\!]+\pi_{\mathsf{GR-RANDOM}}(0,1)$
7:     end for$[\![\gamma[i]]\!]\leftarrow[\![sum]\!]-6$

### v2

@@ 5.3 Secure Evaluation of Synthetic Data
1:Secret shares of binned data [[D]][\![D]\!], Secret shares of binned synthetic data $\hat{D}$]\!], number of patients NN, Domain of gene features $\Omega_{g}={0,1,2,3}$ and classes $\Omega_{y}={0,1,2,3,4}$
2:Secret shares of workload error [[e]][\![e]\!]
3:Initialize [[e]][\![e]\!]
[…]
5:for every value [[x]][\![x]\!] in $rc_{g}]\!]$, $rc_{l}]\!]$, $rc_{l}g]\!]$ do $\triangleright$ Adding noise to guarantee DP
6:  $x]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],(1/N))$
@@
[…]
9:for every value [[x]][\![x]\!] in $\hat{\mu}_{g}]\!]$, $\hat{\mu}_{l}]\!]$, $\hat{\mu}_{l}g]\!]$ do
10:  $x]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],(1/N))$
@@
12:for (every value $\hat{x}$]\!] in $\hat{\mu}_{g}]\!]$, $\hat{\mu}_{l}]\!]$, $\hat{\mu}_{l}g]\!]$) and ([[x]][\![x]\!] in ${\mu}_{g}]\!]$, ${\mu}_{l}]\!]$, ${\mu}_{l}g]\!]$) do
[…]
@@
We note that Line 21 in Algorithm 1 decrements [[c]][\![c]\!] for a data custodian only if both the metrics meet the given thresholds. The data custodians will secret share an array of thresholds for each metric to this end.
@@ 5.4 Empirical Evaluation
[…]
@@
Performance evaluation. We run our framework for different threat models [35, 36, 37] and report the benchmarking runtimes for running the MPC protocols for leukemia data in Table 2. We run for the similar setup as mentioned above, and therefore the input data size $N=945$. For 3PC passive, the parts of the framework with MPC protocol run for a total of $\sim 2.5$ hours – one run of $\pi_{\mathsf{BIN}}$ takes 82.93 secs., $\pi_{\mathsf{INV-BIN}}$ takes 11.83 secs., $\pi_{\mathsf{NOISY-MARG}}$ takes 249.50 secs. $\pi_{\mathsf{LR}}$ with 150 epochs takes 15.42 secs. and $\pi_{\mathsf{WLE}}$ takes 158.66 secs. Table 2 reports performance evalaution for different threat models. The observed runtimes are in-line with the literature and demonstrate that an MPC based framework is feasible for end-to-end synthetic data generation where privacy is of utmost importance compared to the computational time required to generate valuable synthetic data. Future work will focus on optimizing these protocols and proposing new optimized MPC protocols for the proposed framework.
@@ 6 Conclusion
Acknowldgements
This material is based upon work supported by the National Science Foundation under Grant No. 2451163, and by NSF NAIRR 240091 (TACC Frontera). This research was, in part, funded by the National Institutes of Health (NIH) Agreement No. 1OT2OD032581. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the NIH.
@@ Appendix A Secure Multi-Party Computation
Replicated sharing-based 3PC. To give the reader an understanding of how MPC protocols work, we give a brief description of a specific MPC protocol based on replicated secret sharing with three computing parties [35]. A private value $x$ in $\mathbb{Z}_{q}$ is secret shared among servers (parties) $S_{1},S_{2},$ and $S_{3}$ by selecting uniformly random shares $x_{1},x_{2},x_{3}\in\mathbb{Z}_{q}$ such that $x_{1}+x_{2}+x_{3}=x\mod{q}$, and giving $(x_{1},x_{2})$ to $S_{1}$, $(x_{2},x_{3})$ to $S_{2}$, and $(x_{3},x_{1})$ to $S_{3}$. We denote a secret sharing of $x$ by [[x]][\![x]\!].
@@
To multiply secret shared values [[x]][\![x]\!] and [[y]][\![y]\!], we can use the fact that $x\cdot y=(x_{1}+x_{2}+x_{3})(y_{1}+y_{2}+y_{3})\mod{q}$. This means that $S_{1}$ computes $z_{1}=x_{1}\cdot y_{1}+x_{1}\cdot y_{2}+x_{2}\cdot y_{1}\mod{q}$, $S_{2}$ computes $z_{2}=x_{2}\cdot y_{2}+x_{2}\cdot y_{3}+x_{3}\cdot y_{2}\mod{q}$, and $S_{3}$ computes $z_{3}=x_{3}\cdot y_{3}+x_{3}\cdot y_{1}+x_{1}\cdot y_{3}\mod{q}$. The next step is for the servers to obtain an additive secret sharing of 0 by choosing random values $u_{1},u_{2},u_{3}$ such that $u_{1}+u_{2}+u_{3}=0\mod{q}$. This can be done using pseudorandom functions. Each server $S_{i}$ then computes $v_{i}=z_{i}+u_{i}\mod{q}$. Finally, $S_{1}$ sends $v_{1}$ to $S_{3}$, $S_{2}$ sends $v_{2}$ to $S_{1}$, and $S_{3}$ sends $v_{3}$ to $S_{2}$. This allows the servers $S_{1},S_{2}$, and $S_{3}$ to obtain the replicated secret shares $(v_{1},v_{2})$, $(v_{2},v_{3})$, and $(v_{3},v_{1})$, respectively, of the value $v=x\cdot y$.
@@ A.2 MPC Primitives
  • Secure multiplication $\pi_{\mathsf{MUL}}$ : At the start of this protocol, the parties have secret sharings [[a]][\![a]\!] and [[b]][\![b]\!]; at the end, then they have a secret share of $c=a.b$.
  • Secure equality test $\pi_{\mathsf{EQ}}$ : At the start of this protocol, the parties have secret sharings [[a]][\![a]\!]; at the end if $a=0$, then they have a secret share of $1$, else a secret sharing of 0.
  • Secure less than test $\pi_{\mathsf{LT}}$ : At the start of this protocol, the parties have secret sharings [[a]][\![a]\!] and [[b]][\![b]\!] of integers $a$ and $a$; at the end of the protocol they have a secret sharing of $1$ if $a<b$, and a secret sharing of 0 otherwise.
  • Secure less than test $\pi_{\mathsf{ABS}}$ : At the start of this protocol, the parties have secret sharings [[a]][\![a]\!]; at the end of the protocol they have a secret sharing of of its absolute value.
@@ A.3 Protocol for Gaussian Noise
1:Secret shares of drawn sample $\gamma]\!]$
2:Initialize vector $\gamma]\!]$ of length $\max(\omega_{r})$ with 0s
@@
6:   $sum]\!]\leftarrow[\![sum]\!]+\pi_{\mathsf{GR-RANDOM}}(0,1)$
[…]